Grade 12Magical Mathematics[Interesting Approach]

The no. of non negative solutions of the equation:

12x+7y=35 is?

a)0 b)1 c)2 d)3

Plz explain....

Chilukuri Sai Kartik
15 Years agoGrade 12

1 Answer

Veenus Yadav
15 Years ago

by seeing the equation we know that the sum of 12x and 7y must be 35 x any being non negativeconcerning on integral values we get first solution to be (0,5).now if x=1 then y=23/7 and if x=2 then y=11/7 but if x=3 then y=-1/7.thus there are 3 solutions to the eqn

correct option is d